# QADM-Net: Multi-Level Quality-Adaptive Dynamic Network for Reliable Multimodal Classification

### v2

@@ 3.2 Noise-free Prototype Confidence Estimation
3.2.1 Noise-Free Prototype Pre-optimization
For the dataset $\mathcal{D}$ ($M$ modalities and $C$ classes), we aim to find $C\times M$ feature vectors $\omega=\{\{\omega^{c,m}\}_{c=1}^{C}\}_{m=1}^{M}$ termed noise-free prototypes. $\omega^{c,m}$ is a noise-free high confidence representative of the feature vectors of the $m$-th modality from all samples under the $c$-th class. To learn noise-free prototypes, feature vectors of all input samples are first enhanced by modality-specific encoders. Then, the mean of all enhanced feature vectors of different modalities under different classes is calculated and yields $C\times M$ prototypes $\mu=\{\{\mu^{c,m}\}_{c=1}^{C}\}_{m=1}^{M}$. Subsequently, we eliminate noise and enhance the classification confidence of these prototypes $\mu$, optimizing them into $\omega$ by maximizing the constraint $\mathcal{L}^{rob}$: $\omega=\mathop{\arg\max}\limits_{\mu}\mathcal{L}^{rob}$, where $\mathcal{L}^{rob}$ is defined as:
@@ 3.2.1 Noise-Free Prototype Pre-optimization
$cos(\cdot,\cdot)$ is the cosine similarity. $I(\cdot,\cdot)$ refers to mutual information between the two representations. Maximizing the first term of $\mathcal{L}^{rob}$ removes noise and redundant features by increasing the shared information between different modalities Federici et al. (2020). Maximizing the second term improves classification confidence by encouraging orthogonality among different classes Huang et al. (2020).
3.2.2 Multi-level Reliable Quality Estimation
Instead of network structures, a classifier-free design is applied that obtains noise-reliable class probability at both modality and feature levels for each input sample $x_{i}=\{x_{i}^{m}\}_{m=1}^{M}$ by contrasting it with pre-optimized noise-free prototypes $\omega$ at different levels. The modality-level class probability $p_{v}$ is obtained by computing the cosine similarities between the feature vector $x_{i}^{m}$ of each modality $m$ and all noise-free prototypes $\omega^{m}=[\omega^{c,m}]_{c=1}^{C}$ of corresponding modality. For the feature-level probability $p_{f}$, the feature-wise absolute difference between $x_{i}^{m}$ and $\omega^{m}$ is calculated:
@@ 3.2.2 Multi-level Reliable Quality Estimation
After the reliable class probability is obtained, we estimate the modality-level and feature-level quality using the true class probability Han et al. (2022) $\mathcal{T}$: $\delta_{v}(x_{i}^{m})=\mathcal{T}(p_{v}(x_{i}^{m}))\in[0,1],\delta_{f}(x_{i}^{m})=\mathcal{T}(p_{f}(x_{i}^{m}))\in\mathbb{R}^{d^{m}}$. The detailed implementation of $\mathcal{T}$ is given in the supplementary material.
3.3 Global Confidence Normalized Depth
For input sample $x_{i}$, the network depth of QADM-Net is adjusted to $\{\textbf{D}_{i}^{m}\}_{m=1}^{M}$ based on its modality-level qualities $\{\delta_{v}(x_{i}^{m})\}_{m=1}^{M}$. Based on our observation in Figure 1 (a), the lower-quality modality requires a deeper network to enhance its features. Thus, a positive correlation must be maintained between $\tilde{\delta}_{v,i}^{m}=1-\delta_{v}(x_{i}^{m})$ and the network depths $\textbf{D}_{i}^{m}$. Instead of adjusting the depth of each modality for each sample individually, QADM-Net employs a global cross-modality depth normalization strategy. This strategy simultaneously normalizes $\{\{\tilde{\delta}_{v,i}^{m}\}_{m=1}^{M}\}_{i=1}^{N}$ including all modalities from all samples into a unified integer space. The normalized values serve as the corresponding network depths of each modality for each sample $\textbf{D}=\{\{\textbf{D}_{i}^{m}\}_{m=1}^{M}\}_{i=1}^{N}$. On the one hand, this strategy achieves a positive correlation mapping. On the other hand, it offers a simple yet effective way to enhance the reliability of dynamic depth by normalizing the network depth distribution across modalities and mitigating the adverse effects of extreme modality samples. The calculation of this strategy can be formulated as:
@@ 3.3 Global Confidence Normalized Depth
$\mathcal{N}_{K}(\cdot)$ is a function that scales all the input values proportionally to integers within the range $[1,K]$, where $K\geq 1$ is an integer hyperparameter. The analysis of $K$ is provided in the supplementary materials.
@@
Table 1: Comparison with state-of-the-art reliable multimodal classification methods on four datasets.
Method Type BRCA ROSMAP ACC WeightedF1 MacroF1 ACC F1 AUC MD Han et al. (2022) static 87.7$\pm$0.3 88.0$\pm$0.5 84.5$\pm$0.5 84.2$\pm$1.3 84.6$\pm$0.7 91.2$\pm$0.7 MLCLNet Zheng et al. (2023) static 86.4$\pm$1.6 87.8$\pm$1.7 82.6$\pm$1.8 84.4$\pm$1.5 85.2$\pm$1.5 89.3$\pm$1.1 DPNET Zou et al. (2023) static 87.8$\pm$1.0 88.4$\pm$1.2 85.2$\pm$1.2 85.1$\pm$1.1 84.8$\pm$0.7 91.3$\pm$0.7 CALM Zhou et al. (2023) static 88.2$\pm$0.7 88.5$\pm$0.8 85.1$\pm$0.8 85.5$\pm$1.2 87.9$\pm$0.9 91.3$\pm$1.0 QMF Zhang et al. (2023) static 87.4$\pm$0.4 87.7$\pm$0.5 84.1$\pm$0.5 84.6$\pm$0.9 84.8$\pm$0.8 90.5$\pm$0.8 GCFANet Zheng et al. (2024) static 88.6$\pm$1.5 88.9$\pm$1.6 85.3$\pm$1.6 86.3$\pm$1.4 88.3$\pm$1.6 91.5$\pm$1.2 PDF Cao et al. (2024) static 88.2$\pm$0.7 88.0$\pm$0.6 84.9$\pm$0.6 85.9$\pm$0.5 88.0$\pm$0.4 91.6$\pm$0.5 DynMM Xue and Marculescu (2023) dynamic 82.3$\pm$1.7 82.4$\pm$1.7 80.1$\pm$1.2 81.1$\pm$1.5 80.2$\pm$1.6 80.4$\pm$1.3 QADM-Net dynamic 93.4$\pm$ 0.3 93.5$\pm$ 0.4 91.5$\pm$ 0.4 92.2$\pm$ 0.4 92.7$\pm$ 0.3 95.2$\pm$ 0.3 Method Type CUB FOOD101 ACC WeightedF1 MacroF1 ACC WeightedF1 MacroF1 MD Han et al. (2022) static 90.1$\pm$0.7 90.0$\pm$0.8 89.9$\pm$0.7 92.8$\pm$0.3 92.6$\pm$0.2 92.6$\pm$0.2 MLCLNet Zheng et al. (2023) static 88.2$\pm$1.4 88.3$\pm$1.7 87.9$\pm$1.3 92.1$\pm$1.0 92.2$\pm$1.2 92.1$\pm$1.1 DPNET Zou et al. (2023) static 92.1$\pm$0.9 91.8$\pm$0.7 92.2$\pm$1.1 93.1$\pm$1.2 93.0$\pm$0.7 92.8$\pm$0.8 CALM Zhou et al. (2023) static 92.0$\pm$0.2 92.1$\pm$0.4 92.1$\pm$0.5 93.0$\pm$1.1 92.9$\pm$0.7 92.9$\pm$0.8 QMF Zhang et al. (2023) static 89.8$\pm$0.4 89.7$\pm$0.5 89.1$\pm$0.5 92.7$\pm$0.5 92.3$\pm$0.4 92.4$\pm$0.4 GCFANet Zheng et al. (2024) static 92.3$\pm$1.2 92.4$\pm$1.1 92.6$\pm$1.2 92.9$\pm$1.2 93.1$\pm$1.1 92.7$\pm$0.9 PDF Cao et al. (2024) static 93.0$\pm$0.5 93.2$\pm$0.4 93.2$\pm$0.4 93.3$\pm$0.6 93.5$\pm$0.7 93.0$\pm$0.6 DynMM Xue and Marculescu (2023) dynamic 88.3$\pm$1.5 88.3$\pm$1.6 87.7$\pm$1.5 89.3$\pm$1.6 89.4$\pm$1.7 88.5$\pm$1.5 QADM-Net dynamic 97.4$\pm$ 0.2 97.3$\pm$ 0.3 97.4$\pm$ 0.4 94.5$\pm$ 0.3 94.7$\pm$ 0.4 94.4$\pm$ 0.3
3.4 Layer-wise Greedy Parameter
@@
After the depths $\{\textbf{D}_{i}^{m}\}_{m=1}^{M}$ of QADM-Net for the sample $x_{i}$ is determined, the Layer-wise Greedy Parameter (LGP) mechanism predicts the optimal network parameters for all network blocks. A shallow-to-deep layer-wise parameter prediction method is adopted, which can effectively handle network backbones with different depths. At each depth, the greedy strategy predicts the optimal network parameters that maximize the feature-level quality enhancement in the current layer based on how the features have been enhanced in previous depths.
Formally, At depth $t(t\in[1,max(\textbf{D}_{i}^{1},...,\textbf{D}_{i}^{M})])$, we denote the network blocks of all modalities as $B^{t}$, and their inputs of all modalities as $I_{i}^{t}$. Since the depth of some unimodal networks might be less than $t$, $B^{t}$ and $I_{i}^{t}$ can be written as $B^{t}=\{B^{m,t}|t\in[1,\textbf{D}_{i}^{m}]\}_{m=1}^{M}$ and $I_{i}^{t}=Concat(\{\hat{I}_{i}^{m,t}\}_{m=1}^{M})$, where $\hat{I}_{i}^{m,t}$ equals to $I_{i}^{m,t}$ if $t\in[1,\textbf{D}_{i}^{m}]$ and equals to $f_{i}^{m}$ if $t>\textbf{D}_{i}^{m}$. The greedy strategy jointly predicts the parameters $\{\theta_{i}^{m,t}|t\in[1,D^{m}_{i}]\}_{m=1}^{M}$ corresponding to all network blocks in $B^{t}$ utilizing $[I_{i}^{1},...,I_{i}^{t-1}]$, such that the feature-level quality gain $\Delta_{i}^{t}$ at depth $t$ is maximized:
$\Delta_{i}^{t}=mean(max(\delta_{f}(I_{i}^{t+1})-\delta_{f}(I_{i}^{t}),\textbf{0})),$ (7)
where $I_{i}^{t+1}=B^{t}(I_{i}^{t})$. $\delta_{f}$ refers to feature-level quality estimation defined in Section 3.2.2. $max(\cdot,\cdot)$ is the element-wise maximization between two vectors.
@@
In QADM-Net, we use a LSTM unit LU to remember the input feature vectors of previous depths and let $\textbf{LU}^{t-1}$ to have the memory of $[I_{i}^{1},...,I_{i}^{t-1}]$. The input at the $t$-th depth $I_{i}^{t}$ is fed into $\textbf{LU}^{t-1}$ to obtain modality-shared parameter information $\textbf{o}^{t}$ for all blocks of $B^{t}$. Subsequently, the shared information $\textbf{o}^{t}$ is decoded into modality-specific parameters for each block by modality-specific parameter decoders $\{\textbf{PC}^{m}\}_{m=1}^{M}$. The calculation can be formulated as:
$\textbf{o}^{t},\textbf{LU}^{t}=\textbf{LU}^{t-1}(I_{i}^{t}),$
$\theta_{i}^{m,t}=[W_{i}^{m,t},b_{i}^{m,t}]=\textbf{PC}^{m}(\textbf{o}^{t}),\text{if }t\in[1,\textbf{D}_{i}^{m}].$ (8)
@@
Please refer to supplemental material for the pseudo-code of LGP. To guide LGP in predicting the optimal network parameters at each depth, in addition to maximizing $\Delta_{i}^{t}$, a sparsity loss $\mathcal{L}^{s}$ (Equation 9) is introduced to guide the prediction of network parameters that yield the sparsest feature informative vectors $q_{i}^{m,t}$ (mentioned in Equation 2), ensuring maximal removal of irrelevant and redundant information at each layer.
@@
3.5 Learning
During the training stage, we first optimize the noise-free prototypes via maximizing Equation 3. Then, the overall model is trained by minimizing the loss $\mathcal{L}$ defined as:

### v3

@@ III-B Noise-free Prototype Confidence Estimation
III-B1 Noise-Free Prototype Pre-optimization
For the dataset $\mathcal{D}$ ($M$ modalities and $C$ classes), we aim to find $C\times M$ feature vectors $\omega=\{\{\omega^{c,m}\}_{c=1}^{C}\}_{m=1}^{M}$ termed noise-free prototypes. $\omega^{c,m}$ is a noise-free high-confidence representative of the feature vectors of the $m$-th modality from all samples under the $c$-th class. To learn $\omega$, feature vectors of all input samples are first enhanced by modality-specific encoders. Then, the mean of all enhanced feature vectors of different modalities under different classes is calculated and yields $C\times M$ prototypes $\mu=\{\{\mu^{c,m}\}_{c=1}^{C}\}_{m=1}^{M}$. Subsequently, we eliminate noise and enhance the classification confidence of these prototypes $\mu$, optimizing them into $\omega$ by maximizing the constraint $\mathcal{L}^{rob}$: $\omega=\mathop{\arg\max}\limits_{\mu}\mathcal{L}^{rob}$, where $\mathcal{L}^{rob}$ is defined as:
@@ III-B1 Noise-Free Prototype Pre-optimization
$cos(\cdot,\cdot)$ is the cosine similarity. $I(\cdot,\cdot)$ refers to mutual information between the two representations. Maximizing the first term of $\mathcal{L}^{rob}$ removes noise and redundant features by increasing the shared information between modalities [57]. Maximizing the second term improves classification confidence by encouraging orthogonality among classes [58].
III-B2 Multi-level Reliable Quality Estimation
Instead of using network structures or classifiers, a classifier-free design is applied that obtains noise-reliable class probability at both modality and feature levels for each input sample $x_{i}=\{x_{i}^{m}\}_{m=1}^{M}$ by contrasting it with pre-optimized noise-free prototypes $\omega$ at different levels. The modality-level class probability $p_{v}$ is obtained by computing the cosine similarities between the feature vector $x_{i}^{m}$ of each modality $m$ and all noise-free prototypes $\omega^{m}=[\omega^{c,m}]_{c=1}^{C}$ of corresponding modality. For the feature-level probability $p_{f}$, the feature-wise absolute difference between $x_{i}^{m}$ and $\omega^{m}$ is calculated:
@@ III-B2 Multi-level Reliable Quality Estimation
After the reliable class probabilities $p_{v}$ and $p_{f}$ are obtained, the modality-level quality $\delta_{v}$ and feature-level quality $\delta_{f}$ are estimated using the true class probability $\mathcal{T}$ [4]:
$\delta_{v}(x_{i}^{m})=\mathcal{T}(p_{v}(x_{i}^{m}))\in[0,1],$ (6)
$\delta_{f}(x_{i}^{m})=\mathcal{T}(p_{f}(x_{i}^{m}))\in\mathbb{R}^{d^{m}}.$ (7)
The detailed implementation of $\mathcal{T}$ is given in the Supplemental Materials.
Algorithm 1 The layer-wise greedy parameter (LGP) mechanism of Multi-QuAD
Input: Multimodal input samples $\{x_{i}\}_{i=1}^{N}=\{x_{i}^{1},...,x_{i}^{M}\}_{i=1}^{N}$, the depths of Multi-QuAD corresponding to each sample $\{\textbf{D}_{i}\}_{i=1}^{N}=\{\textbf{D}_{i}^{1},...\textbf{D}_{i}^{M}\}_{i=1}^{N}$ provided by the GCND mechanism. The LSTM unit LU. The hidden state dimension number $hdim$. Modality-specific parameter converters $\{\textbf{PD}^{m}\}_{m=1}^{M}$. The feature-level quality estimation function $\delta_{f}(\cdot)$. Random tensor generating function $RT(\cdot)$.
Output: The parameters of each network block of Multi-QuAD for each sample: $\theta_{i}^{m,t}$, reliable classification result of each sample: $\hat{y}_{i}$. The total feature-level quality gain $\Delta$.
1:  $\Delta\leftarrow 0$
2:  for each sample $x_{i}=\{x_{i}^{1},...,x_{i}^{M}\}$ do
3:     Depths of all unimodal networks in Multi-QuAD $\textbf{D}_{i}=\{\textbf{D}_{i}^{1},...,\textbf{D}_{i}^{M}\}$.
4:     The input feature vectors of all modalities at a certain depth: $\{\hat{I}_{i}^{1},...,\hat{I}_{i}^{M}\}$.
5:     The state of the LU at a certain depth: $(h,c)$.
6:     for $t=1$ to $max(\textbf{D}_{i}^{1},...,\textbf{D}_{i}^{M})$ do
7:        if $t=1$ then
8:           $\{\hat{I}_{i}^{1},...,\hat{I}_{i}^{M}\}\leftarrow\{x_{i}^{1},...,x_{i}^{M}\}$.
9:           $(h,c)\leftarrow(RT(hdim),RT(hdim))$.
10:        end if
11:        $I_{i}\leftarrow concat(\hat{I}_{i}^{1},...,\hat{I}_{i}^{M})$.
12:        $\textbf{o}^{t},(h,c)\leftarrow\textbf{LU}(I_{i},(h,c))$.
13:        for $m=1$ to $M$ do
14:           if $t\leq\textbf{D}_{i}^{m}$ then
15:              $\theta_{i}^{m,t}=[W_{i}^{m,t},b_{i}^{m,t}]\leftarrow\textbf{PD}^{m}(\textbf{o}^{t})$
16:              $\hat{I}_{i}^{m+1}\leftarrow\sigma(W_{i}^{m,t}\cdot\hat{I}_{i}^{m}+b_{i}^{m,t})\odot\hat{I}_{i}^{m}$
17:              $\Delta\leftarrow\Delta+mean(max(\delta_{f}(\hat{I}_{i}^{m+1})-\delta_{f}(\hat{I}_{i}^{m}),\textbf{0}))$
18:           end if
19:        end for
20:     end for
21:     The high-quality output feature vectors of sample $x_{i}$: $f_{i}=\{f_{i}^{1},...f_{i}^{M}\}\leftarrow\{\hat{I}_{i}^{1},...,\hat{I}_{i}^{M}\}$.
22:     The reliable classification result of $x_{i}$: $\hat{y}_{i}\leftarrow Classifier(concat(f_{i}^{1},...f_{i}^{M}))$.
23:  end for
III-C Global Confidence Normalized Depth
For input sample $x_{i}$, the network depth of Multi-QuAD is adjusted to $\{\textbf{D}_{i}^{m}\}_{m=1}^{M}$ based on its modality-level qualities $\{\delta_{v}(x_{i}^{m})\}_{m=1}^{M}$. Based on our observation in Fig. 1a, the lower-quality modality requires a deeper unimodal network to enhance its features. Thus, a positive correlation must be maintained between $\tilde{\delta}_{v,i}^{m}=1-\delta_{v}(x_{i}^{m})$ and the network depths $\textbf{D}_{i}^{m}$, where $\delta_{v}(x_{i}^{m})$ is the quality of the modality $m$ input of $x_{i}$ estimated by NFCE. Instead of adjusting the depth of each modality for each sample individually, Multi-QuAD employs a global cross-modality depth normalization strategy. This strategy simultaneously normalizes $\{\{\tilde{\delta}_{v,i}^{m}\}_{m=1}^{M}\}_{i=1}^{N}$ including all modalities from all samples into a unified integer space. The normalized values serve as the corresponding network depths of each modality for each sample $\textbf{D}=\{\{\textbf{D}_{i}^{m}\}_{m=1}^{M}\}_{i=1}^{N}$. On the one hand, this strategy achieves a positive correlation mapping. On the other hand, it offers a simple yet effective way to enhance the reliability of dynamic depth by normalizing the network depth distribution across modalities and mitigating the adverse effects of extreme modality samples. The calculation of this strategy can be formulated as:
@@ III-C Global Confidence Normalized Depth
$\mathcal{N}_{K}(\cdot)$ is a function that scales all the input values proportionally to integers within the range $[1,K]$, where $K\geq 1$ is an integer hyperparameter. The analysis of $K$ is provided in Sec. IV-E3.
@@
TABLE I: Comparison with state-of-the-art reliable multimodal classification methods on four datasets.
Method Type BRCA ROSMAP
ACC WeightedF1 MacroF1 ACC F1 AUC
MD [4] static 87.7$\pm$0.3 88.0$\pm$0.5 84.5$\pm$0.5 84.2$\pm$1.3 84.6$\pm$0.7 91.2$\pm$0.7
MLCLNet [29] static 86.4$\pm$1.6 87.8$\pm$1.7 82.6$\pm$1.8 84.4$\pm$1.5 85.2$\pm$1.5 89.3$\pm$1.1
DPNET [28] static 87.8$\pm$1.0 88.4$\pm$1.2 85.2$\pm$1.2 85.1$\pm$1.1 84.8$\pm$0.7 91.3$\pm$0.7
CALM [3] static 88.2$\pm$0.7 88.5$\pm$0.8 85.1$\pm$0.8 85.5$\pm$1.2 87.9$\pm$0.9 91.3$\pm$1.0
QMF [31] static 87.4$\pm$0.4 87.7$\pm$0.5 84.1$\pm$0.5 84.6$\pm$0.9 84.8$\pm$0.8 90.5$\pm$0.8
GCFANet [1] static 88.6$\pm$1.5 88.9$\pm$1.6 85.3$\pm$1.6 86.3$\pm$1.4 88.3$\pm$1.6 91.5$\pm$1.2
PDF [2] static 88.2$\pm$0.7 88.0$\pm$0.6 84.9$\pm$0.6 85.9$\pm$0.5 88.0$\pm$0.4 91.6$\pm$0.5
DynMM [56] dynamic 82.3$\pm$1.7 82.4$\pm$1.7 80.1$\pm$1.2 81.1$\pm$1.5 80.2$\pm$1.6 80.4$\pm$1.3
Multi-QuAD dynamic 93.4$\pm$ 0.3 93.5$\pm$ 0.4 91.5$\pm$ 0.4 92.2$\pm$ 0.4 92.7$\pm$ 0.3 95.2$\pm$ 0.3
Method Type CUB FOOD101
ACC WeightedF1 MacroF1 ACC WeightedF1 MacroF1
MD [4] static 90.1$\pm$0.7 90.0$\pm$0.8 89.9$\pm$0.7 92.8$\pm$0.3 92.6$\pm$0.2 92.6$\pm$0.2
MLCLNet [29] static 88.2$\pm$1.4 88.3$\pm$1.7 87.9$\pm$1.3 92.1$\pm$1.0 92.2$\pm$1.2 92.1$\pm$1.1
DPNET [28] static 92.1$\pm$0.9 91.8$\pm$0.7 92.2$\pm$1.1 93.1$\pm$1.2 93.0$\pm$0.7 92.8$\pm$0.8
CALM [3] static 92.0$\pm$0.2 92.1$\pm$0.4 92.1$\pm$0.5 93.0$\pm$1.1 92.9$\pm$0.7 92.9$\pm$0.8
QMF [31] static 89.8$\pm$0.4 89.7$\pm$0.5 89.1$\pm$0.5 92.7$\pm$0.5 92.3$\pm$0.4 92.4$\pm$0.4
GCFANet [1] static 92.3$\pm$1.2 92.4$\pm$1.1 92.6$\pm$1.2 92.9$\pm$1.2 93.1$\pm$1.1 92.7$\pm$0.9
PDF [2] static 93.0$\pm$0.5 93.2$\pm$0.4 93.2$\pm$0.4 93.3$\pm$0.6 93.5$\pm$0.7 93.0$\pm$0.6
DynMM [56] dynamic 88.3$\pm$1.5 88.3$\pm$1.6 87.7$\pm$1.5 89.3$\pm$1.6 89.4$\pm$1.7 88.5$\pm$1.5
Multi-QuAD dynamic 97.4$\pm$ 0.2 97.3$\pm$ 0.3 97.4$\pm$ 0.4 94.5$\pm$ 0.3 94.7$\pm$ 0.4 94.4$\pm$ 0.3
III-D Layer-wise Greedy Parameter
@@
After the depths $\{\textbf{D}_{i}^{m}\}_{m=1}^{M}$ of Multi-QuAD for the sample $x_{i}$ is determined, the Layer-wise Greedy Parameter (LGP) mechanism predicts the optimal network parameters for all network blocks in Multi-QuAD. To achieve efficient and reliable parameter prediction on a flexible multimodal framework with dynamic depths, a shallow-to-deep layer-wise greedy prediction strategy is proposed. Specifically, at each depth, the greedy strategy predicts the optimal network parameters jointly across all unimodal networks, aiming to maximize the feature-level quality enhancement in the current layer based on how the features have been enhanced in previous depths. For a certain depth $t\in[1,max(\textbf{D}_{i}^{1},...,\textbf{D}_{i}^{M})]$, denote the “latest” output feature vectors from each unimodal network as $\hat{I}_{i}^{m,t}$:
$\hat{I}_{i}^{m,t}=\begin{cases}I_{i}^{m,t},&\text{if }t\leq\textbf{D}_{i}^{m}\\ f_{i}^{m},&\text{otherwise}.\end{cases}$ (9)
$I_{i}^{m,t}$ is the input feature vector of $B_{i}^{m,t}$ output by $B_{i}^{m,t-1}$, and $f_{i}^{m}$ is the feature vector output by the final layer of $N_{i}^{m}$, as mentioned in Sec. III-A. Let $I_{i}^{t}=Concat(\{\hat{I}_{i}^{m,t}\}_{m=1}^{M})$. The greedy strategy $\textbf{LGP}^{t}$ at the $t$-th layer takes the enhanced feature vectors $I_{i}^{1},\ldots,I_{i}^{t}$ output by the previous $t-1$ layers into account and jointly predicts the parameter values for the $t$-th block of all unimodal networks, such that the feature-level quality gain $\Delta_{i}^{t}$ at depth $t$ is maximized. The mechanism of $\textbf{LGP}^{t}$ at the $t$-th layer can be demonstrated as:
$\{\theta_{i}^{m,t}|t\leq\textbf{D}_{i}^{m}\}_{m=1}^{M}|x_{i}=\textbf{LGP}^{t}(I_{i}^{1},\ldots,I_{i}^{t})$ (10)
[…]
@@
where $\Delta\delta_{f,i}^{m,t}=(max(\delta_{f}(\theta_{i}^{m,t}(I_{i}^{m,t}))-\delta_{f}(I_{i}^{m,t}),\textbf{0}))$. $\delta_{f}(\cdot)$ is the feature-level quality estimation defined in Eq.(7). $N^{m,t}=\sum_{m=1}^{M}{\mathds{1}(t\leq\textbf{D}_{i}^{m})}$. $\theta_{i}^{m,t}(t\leq\textbf{D}_{i}^{m})$ is the parameter value of the $m$-th unimodal network at the $t$-th depth predicted by $\textbf{LGP}^{t}$. As shown in Fig. 3, the detailed implementation of $\textbf{LGP}^{t}$ includes a LSTM unit LU to remember the enhanced feature vectors $I_{i}^{1},\ldots,I_{i}^{t}$ output by previous $t-1$ depths. As $I_{i}^{t}$ input into LU, the output modality-shared parameter information $\textbf{o}^{t}$ is then decoded into modality-specific parameters for the block of each unimodal network by modality-specific parameter decoders $\{\textbf{PD}^{m}\}_{m=1}^{M}$, which can be formulated as:
$\theta_{i}^{m,t}=[W_{i}^{m,t},b_{i}^{m,t}]=\textbf{PD}^{m}(\textbf{o}^{t}),\text{ if }t\leq\textbf{D}_{i}^{m}.$ (11)
@@
The pseudo-code of LGP is provided in Algorithm 1. To guide LGP in predicting the optimal network parameters at each depth, in addition to maximizing $\Delta_{i}^{t}$, a sparsity loss $\mathcal{L}^{s}$ (Eq.(12)) is introduced to guide the prediction of network parameters that yield the sparsest feature informative vectors $q_{i}^{m,t}$ (mentioned in Eq.(2)), ensuring maximal removal of irrelevant and redundant information at each layer.
@@
III-E Learning
During the training stage, we first optimize the noise-free prototypes via maximizing Eq.(3). Then, the overall model is trained by minimizing the loss $\mathcal{L}$ defined as: